OpenStudy (anonymous):

find the four odd consecutive integers whose sum is 132.and find the product of the smallest and largest integer?

OpenStudy (anonymous):

Let x be the first one of the 4. That means that the second one is?

OpenStudy (anonymous):

x+1

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

Not quite. x+1 would be the next integer, but not the next odd one. Each odd integer is how far away from the next odd integer?

OpenStudy (anonymous):

could you please explain me this in detail?

OpenStudy (anonymous):

Lets say that 3 is the first odd integer, What is the next odd integer after 3?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

And 3 is how far from 5

OpenStudy (anonymous):

1

OpenStudy (anonymous):

5-3 = 1?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

I'm pretty sure 5-3 = 2

OpenStudy (anonymous):

right

OpenStudy (anonymous):

So how far away from 5 is the next odd integer?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

sorry 2

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

So if x is the first odd integer in our series, then what are the other 4?

OpenStudy (anonymous):

Err other 3 rather

OpenStudy (anonymous):

X+3,X+5,X+7

OpenStudy (anonymous):

what happen is this right?

OpenStudy (anonymous):

Not quite. The distance between each one is 2. So the first one is x, the second is x+2, the third is x+4, and the fourth?

OpenStudy (anonymous):

x+6

OpenStudy (anonymous):

is this right?

OpenStudy (anonymous):

please give a reply

OpenStudy (anonymous):

Yes. Ok, so now we have the sum of these is 132.

OpenStudy (anonymous):

ok cant we do this like firs intger as x and other integrs like x+1,x+2,x+3,x+4

OpenStudy (anonymous):

No, because if x is 3, the x+1 would be 4 which is not the next odd integer.

OpenStudy (anonymous):

We need them all to be odd, so we need to add 2 every time, not 1

OpenStudy (anonymous):

ok thankyou

OpenStudy (anonymous):

it helped me

OpenStudy (anonymous):

Ok so now you have the equation x + x+2 + x+4 + x+6 = 132. Solve for x.